OpenStudy (anonymous):

Find the exact solutions of sin^2X + Sin2X = -Cos^2X + 1 in the interval [0, 2pi)

myininaya (myininaya):

Hey did you try putting everything on one side

myininaya (myininaya):

And then us \[\cos^2(x)=1-\sin^2(x)\]

myininaya (myininaya):

use*

OpenStudy (anonymous):

No.... I didn't try that. But I have no idea how to do it. can you take it from the top...

myininaya (myininaya):

Put everything on one side

myininaya (myininaya):

add cos^2(x) on both sides subtract 1 on both sides

OpenStudy (anonymous):

yeh, I have it all on 1 side.

myininaya (myininaya):

ok the use that identity I mentioned and replace cos^2(x) with 1-sin^2(x) tell me if anything cancels :)

OpenStudy (anonymous):

Got it

OpenStudy (anonymous):

I got the pi/2 and 3pi/2 but there's two more answers (0 and pi) which I don't get how they got.

OpenStudy (anonymous):

you just have to write the equation like that sin^2(x)+cos^2(x)+sin(2x)=1

OpenStudy (anonymous):

I got to sin2X=0 then solved for x= pi/2 +2 n pi and X=pi + 2 n pi

myininaya (myininaya):

\[0 \le x < 2 \pi\] \[\sin(2x)=0\] if we let u=2x then we have \[\sin(u)=0 \text{ where } 0 \le \frac{u}{2} < 2 \pi \] \[\sin(u)=0 \text{ where} 0 \le u < 4 \pi\] What are values of u that satisfy this equation between 0 and 4 pi include 0

OpenStudy (anonymous):

Oh... Okay. How are you making the sign for pi and x? (For future reference)

myininaya (myininaya):